Crew 215 Commander Report 20Nov2019

[title Commander Report – November 20th]

Crew 216 Commander Report 26-NOV-2019
Sol: 2
Summary Title: Getting into routine
Author’s name: Marc Levesque
Crew 216 started our second sol on the Red Planet as all members continue to settle in and establish working routines for life in isolation and confinement. Everyone exhibits a very positive attitude for helping each other out in daily tasks. I am also imressed by everyone’s upbeat nature and ability to roll with any hiccups they might encounter. I feel very fortunate to have been blessed with an extremely talented crew and look forward to all maintaining an upbeat attitude.

[end]

Crew 216 Operations Report 26Nov2019

[title Operations Report – November 26th]

Crew 216 Operations Report 26-Nov-19
SOL: 2
Name of person filing report: Mike Lawson
Non-Nominal systems:
Notes on Non-Nominal systems:
Generator: run
Hours run: 12.3
From what time last night: 1900
To what time this morning: 0720
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night: 75% as of 1620
Diesel Reading – 42
Station Propane Reading – 60
EthaNol Free Gasoline: N/A
Water (loft tank) (gal): 47
Water Meter (units): 1461497
Water (static tank) (gal): 421
Static to Loft Pump used – Yes
Water in Green Hab (gal): 274
Water in Science Dome (gal): 0
Toilet tank emptied: No
Deimos rover used: No
Hours:
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used:
Hours:
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used: Yes x2 EVAs
Hours: 114
Beginning charge: (Before EVA): 100
Ending charge: (On return from EVA, before recharging): 48
Currently charging: Yes
Opportunity rover used: Yes
Hours: 68.7
Beginning charge: 100
Ending charge: 46
Currently charging: Yes
Curiosity rover used: Yes
Hours: 119.4
Beginning charge: 100
Ending charge: (On return from EVA, before charging):70
Currently charging: Yes
Notes on rovers: 0
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No
Reason for use: N/A
Oil Added? No
# Hours the ATVs were used today: 0
Notes on ATVs:
HabCar used and why, where?
CrewCar used and why, where?
General Notes and comments:
Summary of internet: No serious issues.
Summary of suits and radios: Open VOX on radio in HAB caused issues with morning EVA. Resolved.
Summary of Hab operations: Routine maintenance and housekeeping.
Summary of GreenHab operations: Routine watering.
Summary of Science Dome operations: Gypsum samples processing started.
Summary of RAM operations: No issues.
Summary of any observatory issues: Visual observation of sun.
Summary of health and safety issues: Icing on rear HAB airlock, removed by engineer. CO test negative.
Questions, concerns and requests to Mission Support: None

Crew 216 Sol Summary 26Nov2019

[title Sol Summary – November 20th]

Crew 216 Sol Summary Report 26-NOV-19
Sol: 2
Summary Title: Communications, film, and drone projects
Author’s name: Marc Levesque
Mission Status: Active
Sol Activity Summary: Today was our second day in sim during which we conducted two EVAs. The first was to test an improved communication system for EVAs using a VHF/UHF 5 watt radio at the Hab and to continue filming for a documentary project. The second continued communications testing to determine the range of the GMRS radio at the Hab versus the VHF/UHF radio and to conduct the first test of the drone project.
Look Ahead Plan: We will conduct two EVAs to identify possible radio repeater sites near the Hab and to continue testing EVA communications to the south of the Hab along with documentary filming.
Anomalies in work: N/A.
Weather: Cold and windy overnight.
Crew Physical Status: Healthy
EVA: Two EVAs conducted.
Reports to be file: Sol Summary Report, Commander Report, Operations Report, Astronomy Report, GreenHab Report, Journalist Report, Research Report, EVA Report, and Photos.
Support Requested: None

Crew 216 Science Report 26Nov2019

[title Science Report – November 26th]

Crew 216 Research Report 26-NOV-19
Sol 2
Submitted by: Rich Whittle

1) Psych Scale (TAMU IRB2019-1042D) – Nothing further to report.

2) Medical Casts – Continuation of Crew 215 research. Crew washed gypsum and experimented with crushing techniques

3) In-Situ Fluorescent Mineral Prospecting – Still searching for radioactives and rare Earths.

4) MDRS Mapping – Nothing further to report (awaiting return of map with desired annotations).

5) MDRS Communications – Attempted to test increased range of 5W transceiver to south, as well as find potential rebroadcast locations. EVA cut short due to Hab comms issue, though one potential site was marked at grid 12 S 0518585 4248798. Repeat attempt planned for EVA 06 on 27-Nov-19. On EVA 04 traveled to east of Hab to evaluate comms. Good comms as far as crew traveled, distance cut short due to battery life of rovers. EVA 05 planned tomorrow to evaluate radio rebroadcast sites close to Hab.

[end]

Crew 216 Journalist Report 26Nov2019

[title Journalist Report – November 26th]

Crew 216 Journalist Report 26-NOV-19
Sol 2
Author’sName: Evgenia Alexandrova
Today’s report is all about numbers. It’s Sol 2 on Mars and many things happened for the first time today. Two crew members accomplished their first EVA, while another rinsed seven handfuls of the gypsum collected the day before. There was a first Martian shower (quantity of hot water needed – 3 boiled kettles), calculation of the airlock volume (175.94 cubic feet), and 3,000 words written for the new paper. Ending report with a new poem, this time only numbers (to be read out loud):
1 11 25
25 11
6 13 and 1025
7 7 7

Crew 216 Greenhab Report 26Nov2019

[title Greenhab Report – November 26th]

Crew 216 GreenHab Report 26-NOV-19
GreenHab Officer: Michael Ho
Environmental control: Door left closed for the day.
Average temperature: 23C
Hours of supplemental light: Automatic system
Daily water usage for crops: 16L ( blue cans)…of which 6L were from a 3/4-full watering can left from previous rotation.
Daily water usage for research and/or other purposes: nil
Water in Blue Tank – enough
Time(s) of watering for crops: 10:00 & 16:30
Newly sprouted:
from 1 pot of Swiss Chard,
from 1 pot of tomatoes,
from 1 pot of pepper colour mix,
from both long trays of nassissus,
from 1 round tray of snow peas.

SEASONAL PLANTING HISTORY:
1 more avocado seed planted today.
3 terracotta pots of 4 succulents each planted today, placed on the same shelf as the aloe vera. Soil mix: Potting soil and local sand.

(Established plants introduced since 03/11/2019 till 22/11/2019 please refer to Crew 215 GreenHab Report of 22/11/2019.)

OTHER NOTES
Harvest: None
Support/supplies requested: None

Crew 216 EVA Report 26Nov2019

Crew 216 EVA Report 26-Nov-2019
EVA # 04
Author: Kennedy
Purpose of EVA: Evaluate communication range to east of Hab with 5W radio transceiver, explore locations for communications rebroadcast, collect drone footage for education.
Start time: 1400
End time: 1545
Narrative: We began with a small technical difficulty with the drone footage. Michael had to replace the battery and reevaluate the auto follow feature. He eventually manually operated the drone while we drove the vehicles. Once in transit the drive to Galileo Road was uneventful. Near the junction of Galileo Rd 1104 and Cactus Rd 1104 the road travels in the wash. Due to recent precipitation, signs of vehicle travel were erased, we had to scout on foot to confirm the route that was in the GPS and maps was accurate. After a short survey of the area we found the rest of the road and continued down Cactus Rd 1104. At 0520999E 4252250N we stopped and evaluated the area. Comms were solid as there are few obstructions between that area and the Hab. The area was still quite grassy and not terribly Martian. We returned without incident.
Destination: Eastern end of Cactus Rd 1104
Coordinates (use UTM NAD27 CONUS): 12 S 0521940 4252800.
EVA Participants: Kennedy, Ho, Levesque
Road(s) and routes per MDRS Map:
1) Hab to 12 S 0519100 4251980 via Cow Dung Rd,
2) 12 S 0519100 4251980 to 12 S 0519880 4251540 via Galileo Rd 1104
3) 12 S 0519880 4251540 to east end of Cactus Rd 1104
4) Explore suitable areas either side of road on foot
5) Return to Hab via reverse route
Mode of travel: Rover/Walk
Vehicles used: Opportunity, Spirit

[end]

Crew 216 EVA Report 26Nov2019

[title EVA Report – November 26th]

Crew 216 EVA Report 26-Nov-2019
EVA # 03
Author: Whittle
Purpose of EVA: Evaluate communication range to S of Hab with 5W radio transceiver, explore locations for communications rebroadcast, collect more footage for documentary.
Start time: 0900
End time: 0945
Narrative: Crew left Hab to test range of 5W radio transceiver. Continued radio checks every minute until lost comms with Hab when passing through Kissing Camel Ridge (this had been the hypothesized loss of comms point due to terrain) at approximately 0905. Crew parked on road at grid 12 S 0518682 42488798. After discussing plan to evaluate high points at 12 S 0518585 4248798 and 12 S 0518758 4249076 (West and East of Cow Dung Rd), crew climbed West point (12 S 0518585 4248798). Here crew reestablished communications with Hab at 0915, discussed plan to evaluate second point to East of Cow Dung Rd (12 S 0518758 4249076). Crew captured landscape footage here. On route to 12 S 0518758 4249076 at approximately 0930 crew received message from Hab instructing return to Hab due to comms issue. Crew complied and returned to Hab, arriving at 0940. It should be noted for the report that crew maintained good internal EVA communications at all times (save for approximately 2 minutes after exiting rovers when ARES 9’s microphone fell off his head), and had good communications with Hab at all times except for the aforementioned loss of communications between passing through Kissing Camel Ridge and arriving at the high point at grid 12 S 0518585 4248798. On return, crew was made aware of communications issues at Hab, these did not affect internal comms amongst EVA members on channel 22, or communication from EVA to Hab using the 5W transceiver.
Destination: High Point South of Kissing Camel Ridge
Coordinates (use UTM NAD27 CONUS): 12 S 0518585 4248798.
Participants: Whittle, Lawson, Alexandrova
Road(s) and routes per MDRS Map:
1) Hab to 12 S 0518682 4248879 via Cow Dung Rd
2) Walked from 12 S 0518682 4248879 to high point at 12 S 0518585 4248798
3) Walked from 12 S 0518585 4248798 to 12 S 0518682 4248879
4) Returned to Hab via Cow Dung Rd
Mode of travel: Curiosity, Spirit

Astronomy Report- November 26th

Crew 216 Astronomy Report 26-NOV-19
Name: Michael Ho

MDRS ROBOTIC OBSERVATORY

Robotic Telescope Requested: none

Objects to be Imaged this Evening: NGC 281

Images submitted with this report: none

Problems Encountered: none

MUSK OBSERVATORY

Solar Features Observed: Good sunny morning, getting the hang of observation. No prominences observed.

Images submitted with this report: none

Problems Encountered: Waiting for cable for astronomy laptop for imaging by camera.